The rate of flow, Q, of liquid in a cylindrical tube is given by the equation

Q = πr4 ΔP/(8ηL), where

Q = rate of flow of liquid (m3/s)

r = radius of the tube, (m)

L = length of the tube (m)

ΔP = pressure difference between the ends of the tube (Pa)


i. Find the dimensions of the dimensions of the quantities r4 , Q and ΔP.

ii. Determine the dimensions of the quantity η.



1
Expert's answer
2021-05-31T15:12:55-0400

Rate of flow, Q, of liquid in cylindrical tube is given by the equation -


"Q=\dfrac{{\pi}r^{4}{\Delta P}}{8{\eta}L}"


Q = rate of flow of liquid "(m^{3}\/s)"

r = radius of tube (m)

L = length of tube in (m)

"{\Delta}P=" pressure between ends of the tube (pa).


dimensions of all the term given in formula -


"r=L"

P "=ML^{-1}T^{-2}"

"{\eta}=ML^{-1}T^{-1}"


1)

Now putting these dimensions in above formula , we get -


"Q=\dfrac{[Ml^{-1}T^{-2}][L^{4}]}{[ML^{-1}T^{-1}][L]}"


"Q=L^{3}T^{-1}"


Now finding the dimensions of "r^{4}" -


"r^{4}=\dfrac{{\eta}LQ}{{\Delta P}}"


Now putting the dimensions in above , we get -



= "r^{4}=\dfrac{[ML^{-1}T^{-1}][L][L^{3}T^{-1}]}{[ML^{-1}T^{-2}]}"


"=r^{4}=L^{4}"


"\Delta{P}=\dfrac{{\eta}LQ}{{r^{4}}}"


now putting the dimensions , in above equation we get -



"{\Delta}P=\dfrac{[ML^{-1}T^{-1}][L][L^{3}T^{-1}]}{L^{4}}=" "ML^{-1}T^{-2}"


2)


Dimensions of quantity "{\eta}=\dfrac{r^{4}{\Delta P}}{LQ}"


Now putting the dimensions the quantity , in the above question ,we get ,


"{\eta}=\dfrac{[L^{4}][ML^{-1}T^{-2}]}{[L][L^{3}T^{-1}]}" "=" "ML^{-1}T^{-1}"